• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name f.cusanno

    Log entry time 07:40:28 on December10,2001

    Entry number 79304

    keyword=shift summary

    12/10 OWL shift
    Shift crew: F.Cusanno, F.Garibaldi

    Taking data for the RICH tests.
    Left_arm = hadron
    Right_arm = electrons

    00:00 working with the trigger, daq and pedestal runs
    Found a stable setting :
    RICH HV = 2150 V
    Vgrid = 300 Volts
    delay = 396ns
    Trigger with S0
    Pedestals: new 'pedestal.dat' with 4 sigma threshold
    Current = 20 uA
    Target : Carbon thick

    Run 3081 -- 3083 are pedestal runs
    Run 3084 -- 3086 are junk
    Run 3087 are a test run

    01:36 START the delay scan with run 3096
    03:04 MCC called: they do QE measurements
    03:07 Beam is back

    RunNr Trigger Type

    3088 T3 and T5, no S0
    3089 T3 and T5, no S0
    3090 Junk
    3091 Junk
    3092 Junk
    3093 T3 and T5, no S0
    3094 Junk
    3095 S0 and T5, no T3


    RunNr........delaySettings (+244ns)......NrEvents........Q(uA*hours)
    3096...........8 48 40 48 8 = 152 .............350K...............9.1
    3097...........104 48 40 48 104 = 344 .....500K...............13.6
    3098...........56 48 40 48 104 = 296........500K...............13.7
    3099...........56 48 40 48 56 = 248........500K...............13.9
    3100...........56 48 40 48 8 = 200........500K...............13.7
    3101........... 8 48 40 48 8 = 152..........1M ................27.4
    3102............8 48 40 48 8 =152.........130K............... 3.6

    End of the tests